OpenStudy (anonymous):

Help please ! a country's population in 1992 was 39 million. In 1997 it was 42 million. Estimate the population in 2009 using the exponential growth formula

OpenStudy (anonymous):

Xt = Xo (1+r)^t Xt = population at time t Xo = population at time 0 t = time r = rate. At 1992, 39 = Xo (1+r)^1992 ------(1) At 1997 42 = Xo (1+r)^1997 -------- (2) (1) / (2): 39/42 = (1+r)^(1992-1997) 39/42 = (1+r)^(-5) 39/42 = 1 / (1+r)^5 (1+r)^5 = 42 / 39 (1+r) = root5 (42/39) r = root5(42/39) -1 r = 0.014931978 Thus, 42 = Xo (1+0.014931978)^1997 Xo = 5.870657669 x 10^-12 At 2009, X2009 = (5.870657669X10^-12) (1+0.014931978)^2009 X2009 = 50.17669491 MILLION

OpenStudy (anonymous):

thanks !!!!